Problem 17.75 A solution contains three anions witIf a dilute AgNO3 solution is slowly added to the solution, what is the first compound to precipitate: Ag2CrO4 ( Ksp = 1.2×10?12), Ag2CO3 ( Ksp = 8.1×10?12), or AgCl ( Ksp = 1.8×10?10)?
Answer #1

You have to apply the definition of Ksp (constant equilibrium solubility):

for the first one CrO4-2

Ag2CrO4 === 2 Ag+ + CrO4-2

........................2x.....................x

we already know the concentration of [CrO4] = 0.2

ksp = 1.2 x 10-12 = 4x2 * 0.2

4x2 = 1.2 x 10-12 / 0.2

x = 6.12 x 10-7 M

for the second ion

Ag2CO3 ==== 2 Ag + CO3

.........................2x..............x

ksp = 8.1 x 10-12

ksp = (2x)2 * x

concentration of CO3 is 0.1M

8.1 x 10-12 = 4x2 * 0.1

x = 4.5 x 10-6

fo the chloride

AgCl ==== Ag + Cl

...................x.........x

ksp = 1.8 x 10-10 = x (0.01)

the concentration of cl is 0.01

x = 1.8 x 10-8 M

according to the calculations we need a small amount of Cl (compared to the other anions) to produce a precipitate so the first compound to precipitate is AgCl

*hope it helps =)